# Littlewood-Offord bounds on the symmetric groups and applications

### v1

@@ Abstract.
The paper contains several new results, addressing both discrete and continuous anti-concentration phenomena. On the discrete side, we establish a near-optimal structural characterization of the vectors $\bm{w}$} and $\bm{v}$} under the assumption that the concentration probability $\sup_{x}{\mathbb{P}}(S_{\pi}=x)$ is polynomially large. As applications, we derive and strengthen a number of previous results. In particular, we show that if both $\bm{w}$} and $\bm{v}$} have distinct entries, then $\sup_{x}{\mathbb{P}}(S_{\pi}=x)\leq n^{-5/2+o(1)}.$ This bound serves as a permutation-space analogue of the classical Erdős–Moser bound in the product-space setting and answers a question posed by Alon–Pohoata–Zhu [2].
From the continuous perspective, we study the small-ball event $|S_{\pi}-L|\leq\delta.$ We establish sharp bounds in various settings, including results exhibiting sub-gaussian decay in $L$, thereby settling a question of Söze [50]. With additional effort, we are also able to treat the joint distribution of these events. Moreover, we provide a characterization of the vectors $\bm{w}$} and $\bm{v}$} for which these small-ball probabilities are large. As an application, we prove that the number of extremal points of random permutation polynomials is bounded by $O(\log n)$, extending results of Söze [49, 50] on the number of real roots.
H. Nguyen is supported by a Simons Travel Grant TSM-00013318. T. Tran is supported by the Excellent Young Talents Program (Overseas) of the National Natural Science Foundation of China. A substantial part of this work was carried out during the ADM program 2024; the authors are grateful to the VinBigData Center for its hospitality.
@@ 1.1. Anti-concentration in product spaces
$\sup_{x}{\mathbb{P}}(S=x)=O\!\left(\frac{1}{\sqrt{n}}\right).$
Here and throughout, asymptotic notation is taken in the limit $n\to\infty$.
@@
Let $W$ denotes the multi-set $\{w_{1},\dots,w_{n}\}$. Let
@@
Assume that $\rho(W)\geq n^{-C}$ for some constant $C>0$. This implies that at least $2^{n}n^{-C}$ of the $2^{n}$ subsums of $W$ coincide, suggesting that $W$ must possess substantial additive structure. Tao and the last author managed to formalize this intuition in a quantitative way via the notion of generalized arithmetic progressions (GAPs).
@@
[39, Theorem 2.1] Let $C>0$ and ${\varepsilon}\in(0,1)$ be constants. Suppose $W=\{w_{1},\dots,w_{n}\}$ is a multiset of real numbers such that
@@
There are many extensions to other frameworks, such as the extension from linear forms $S$ of the $\xi_{i}$ to quadratic and higher-order multilinear forms by Costello, Nguyen [10, 36], and Meka et al. [35], and to non-abelian groups by Tiep–Vu [43] and Juskevicius–Semetulskis [26] (see also [29, 38, 54]). Other notable contributions include the works of Tao–Vu [57], Rudelson–Vershynin [47], and, more recently, Fox et al. [15] and Kwan et al. [30, 31]. The discussion of these papers are beyond the scope of our works.
The continuous case proceeds in parallel with the discrete case, and we refer the reader to the survey [40] for further discussion. In particular, the continuous analogue of Theorem 1.6 asserts that the elements of $W$ lie close to a small GAP.
1.2. The space of random permutations: discrete setting
@@
1.2.1. Some forward theorems
In this subsection, we bound the probability that $S_{\pi}=x$ for any fixed value $x$. This problem has been studied recently in the combinatorics community. In [50], Söze considered the special case ${\bm{v}}=(1,2,\dots,n)$ and proved the following result.
@@ 1.2.1. Some forward theorems
$\sup_{x}{\mathbb{P}}(S_{\pi}=x)=O\!\left(\frac{1}{n}\right).$
@@
$\sup_{x}{\mathbb{P}}(S_{\pi}=x)=O\!\left(\frac{1}{n}\right).$
Both Theorems 1.7 and 1.8 are special cases of continuous theorems that we will discuss later. Theorem 1.8 may be viewed as a permutation analogue of the Littlewood–Offord–Erdős result from the product-space setting. Note that there is a $\sqrt{n}$ improvement from the bound $O(n^{-1/2})$ to $O(n^{-1})$.
@@
The proof of Theorem 1.9 is quite involved and relies on tools from algebraic combinatorics. Roughly speaking, if $\bm{v}$} has distinct coordinates and $S_{\pi}=x$, then the set of such permutations $\pi$ forms an antichain (in a suitable weakening of the Bruhat order on $S_{n}$). This antichain can be identified with a disjoint union of Bruhat orders on $S_{n}/S_{\alpha}$ for certain parabolic subgroups $S_{\alpha}$. These posets have the Sperner property, which implies that their size is bounded by the largest rank of $S_{n}/S_{\alpha}$. These ranks correspond to coefficients of the $q$-multinomial coefficient $\binom{n}{\alpha}_{q}$, and Theorem 1.9 follows from appropriate bounds on those coefficients.
While Theorem 1.9 is elegant—being optimal and valid for all nn—the above methods do not appear to extend easily to more general settings where additional structure is imposed on $\bm{v}$} and $\bm{w}$}. As observed by Alon et al. in [2], if both the $v_{i}$ and the $w_{i}$ are distinct, then ${\mathbb{P}}(S_{\pi}=x)$ must lie between $n^{-3/2}$ and $n^{-5/2}$. A sharp bound in this regime would serve as a permutation-space analogue of the Erdős–Moser result in the product-space setting. However, the techniques of [2] and  [42] do not resolve this problem. We will address it in Section 1.4.
1.2.2. An inverse theorem
One of the main results of our paper is the following analogue of Theorem 1.6, providing a characterization of those vectors $\bm{w}$} and $\bm{v}$} for which $S_{\pi}$ admits polynomially large concentration.
@@ 1.2.2. An inverse theorem
Then there exists a proper symmetric GAP $Q$ of rank $r=O_{C,{\varepsilon}}(1)$ and size $O_{{\varepsilon}}(\rho^{-1}n^{-r/2})$ that contains $(w_{i}-w_{j})(v_{k}-v_{l})$ for at least $(1-{\varepsilon})n^{4}$ quadruples $(i,j,k,l)$.
The bound is sharp; we refer the reader to Lemma 3.2 for a proof. In Section 1.4, we will use this theorem to derive and refine various forward results. In particular, we will resolve the question posed by Alon et al. discussed in the previous subsection.
1.2.3. A more general setting with two dimensional arrays
@@
Then there exists a proper symmetric GAP $Q$ of rank $r=O_{C,{\varepsilon}}(1)$ and size $O_{{\varepsilon}}(\rho^{-1}n^{-r/2})$ that contains $a_{ik}-a_{jk}-a_{il}+a_{jl}$ for at least $(1-{\varepsilon})n^{4}$ quadruples $(i,j,k,l)$.
We also refer the reader to Remark 3.1 for a more detailed discussion of the implications of these structural conclusions. Theorem 1.10 corresponds to the special case $a_{ij}=w_{i}v_{j}$.
1.3. The space of random permutations: continuous setting
In the continuous setting, we study events of the form $S_{\pi}\in I$ for a small interval $I$. In this context, it is more natural to assume that $\|{\bm{v}}\|_{\infty}=O(1)$. We will often write such events as
$|S_{\pi}-L|\leq\delta,$
where $L$ is the center of the interval and $\delta$ its radius. In probability theory, bounds for events of this type are commonly referred to as small-ball probabilities. It is clear that the continuous problem is strictly harder, as it contains the discrete problem as a special case (one can shrink the interval to a single point).
In the product-space setting, this problem has been studied extensively, with early foundational results due to Rogozin [46], Kolmogorov [28], and Halász [17]. There have also been many recent developments, especially in the inverse direction; see, for instance, [39, 47, 58] and the survey [40].
For $S_{\pi}$, CLT results and Berry–Esseen–type estimates have been investigated beginning with the classical works of Wald–Wolfowitz [41, 62] and Hoeffding [18] in the early 1950s. More recent approaches based on the Lindeberg exchange method or Stein’s method can be found in [1, 6, 7, 8, 44]. We refer the reader to [9] and the references therein for further contributions concerning this important statistic. The best anti-concentration bound obtainable from these classical approaches is of order $O(n^{-1/2})$ (independent of the length of the interval), arising from the rate of convergence in the CLT.
Although our methods have the potential to work under very general conditions on $\bm{w}$} and $\bm{v}$} (see, for instance, Theorem 1.22), in what follows we restrict our attention to one of the most natural choices for $\bm{v}$}, namely sequences arising from polynomials.
@@ 1.3.1. A starting point: the linear case $v_{i}=i/n$
Our starting point is the following result of Söze [50, Lemma 4], in which he studied $S_{\pi}$ from (2) for the special choice of $\bm{v}$} given by $v_{i}=i/n$.
@@
An interesting feature of this theorem is the appearance of the parameter $L$ in the bound, which implies that $S_{\pi}$ exhibits exponential decay. Söze [50] conjectured a sharper estimate of order $O\!\bigl(n^{-1}e^{-\Theta(L^{2})}\bigr)$, corresponding to sub-gaussian decay. Notice that once $L$ appears on the right-hand side, the magnitude of the $v_{i}$ becomes relevant.
@@
Other natural questions include weakening the strong restriction ${\bm{v}}=(1,\dots,n)/n$ and treating scales smaller than $1/n$. We will achieve these goals under a modest assumption on $\bm{w}$}. Namely, throughout what follows, we assume the following.
Condition 1.14 (Non-degeneracy).
@@
Roughly speaking, this condition corresponds to $\bm{w}$} being a unit vector orthogonal to $\mathbf{1}$ and satisfying $\|{\bm{w}}\|_{\infty}\ll 1/\sqrt{\log n}$.
@@
Let $\delta>0$ be given. Suppose that the sequence $(w_{1},\ldots,w_{n})$ satisfies Condition 1.14 for some sufficiently large constant $A>0$. Let $I\subset[n]$ be any subset with $|I|\geq\delta n$, and consider a sequence $(v_{1},\ldots,v_{n})$ that is partially specified by
@@
${\mathbb{P}}\Bigl(\bigl|S_{\pi}-L\bigr|\leq\frac{1}{n}\Bigr)=O_{A}\!\left(\frac{1}{n}\right).$ (5)
@@
${\mathbb{P}}\Bigl(\bigl|S_{\pi}-L\bigr|\leq\frac{1}{n}\Bigr)=O\!\left(\frac{1}{n}e^{-\Theta(L^{2})}\right).$ (6)
@@
The answer to Problem 1.13 is affirmative for any vector $\bm{w}$} satisfying Condition 1.14.
@@
Let $0<\varepsilon<1/2$ and $\delta>0$ be given. Suppose that the sequence $(w_{1},\ldots,w_{n})$ satisfies Condition 1.14 for some sufficiently large constant $A>0$. Furthermore, assume that no interval of length $\varepsilon/\sqrt{n}$ contains more than $(1-\varepsilon)n$ of the values $w_{i}$. Let $I\subset[n]$ be any subset with $|I|\geq\delta n$, and consider a sequence $(v_{1},\ldots,v_{n})$ that is partially specified by
@@
${\mathbb{P}}\Bigl(\bigl|S_{\pi}-L\bigr|\leq\frac{1}{n^{3/2}}\Bigr)=O_{A}\!\left(\frac{1}{n^{3/2}}\right).$ (7)
@@
${\mathbb{P}}\Bigl(\bigl|S_{\pi}-L\bigr|\leq\frac{1}{n^{3/2-\varepsilon}}\Bigr)=O\!\left(\frac{1}{n^{3/2-\varepsilon}}e^{-\Theta(L^{2})}\right).$ (8)
@@
This result may be viewed as a continuous analogue of Theorem 1.23. We emphasize that the condition excluding intervals of length $1/\sqrt{n}$ that contain most of the $w_{i}$ is essential for obtaining small-ball estimates at the $n^{-3/2}$ scale. In the $L$-dependent bound (8), we assume the slightly larger radius $n^{-3/2+\varepsilon}$ purely for technical convenience; see our treatment of the “very large |t||t|” regime in the proof of (8) in Section 4.
@@
Finally, we remark that Theorem 1.15 may still hold without Condition 1.14. However, for Theorems 1.17 and 4.4, additional assumptions (of the type imposed above) are necessary, since these results fail dramatically when only a few of the $w_{i}$ are nonzero.

### v2

# The anti-concentration phenomenon with respect to random permutations

@@ Abstract.
The paper contains several new results, addressing both discrete and continuous anti-concentration phenomena. On the discrete side, we establish a near-optimal structural characterization of the vectors $\bm{w}$} and $\bm{v}$} under the assumption that the concentration probability $\sup_{x}{\mathbb{P}}(S_{\pi}=x)$ is polynomially large. This is an “inverse theorem” and can be seen as an analog of earlier results by Tao–Vu and Nguyen–Vu in the product space setting. In fact, our technique allows to obtain an intervese theorem for random subsums of entries of a matrix, which contains the result concerning random permutation as the special case where the matrix has rank one.
On the continuous side, we study the small-ball event $|S_{\pi}-L|\leq\delta$. Our results exhibit sub-gaussian decay in $L$, answering an open question of Söze [54]. With additional effort, we are also able to treat the joint distribution of these events. We also extend the inverse theorem obtained in the discrete case to the continuous case.
Our results have applications in various areas. First, we use our inverse theorems to derive and strengthen a number of previous anti-concentration bounds. In particular, we show that if both $\bm{w}$} and $\bm{v}$} have distinct entries, then $\sup_{x}{\mathbb{P}}(S_{\pi}=x)\leq n^{-5/2+o(1)}.$ This bound serves as an analog of the classical Erdős–Moser bound in the product-space setting and answers a question posed by Alon–Pohoata–Zhu [2]. Next, we apply our new results to study random polynomials, and prove that the number of extremal points of random permutation polynomials is bounded by $O(\log n)$, extending results of Söze [53, 54] on the number of real roots. In the final application, we prove that random matrices whose rows are independent random permutations of a fixed non-degenerate vector are nonsingular with high probability, strengthening earlier results of Nguyen and Nguyen–Vu [40, 44].
H. Nguyen is supported by a Simons Travel Grant TSM-00013318. T. Tran is supported by the Excellent Young Talents Program (Overseas) of the National Natural Science Foundation of China GG0010007003. A substantial part of this work was carried out during the ADM program 2024 supported by VinIF foundation; the authors are grateful to the VinUni BigData Research Insitute for its hospitality.
@@ 1.1. Anti-concentration in product spaces
$\sup_{x}{\mathbb{P}}(S=x)=O\Big(\frac{1}{\sqrt{n}}\Big).$
Notation. Here and throughout, asymptotic notation is taken in the limit $n\to\infty$. For real quantities $X,Y$, we write $X=O(Y)$, or equivalently $Y=\Omega(X)$, if $|X|\leq C|Y|$ for some constant $C>0$ independent of $n$. This constant $C$ may depend on other fixed quantities. If $X=O(Y)$ and $Y=O(X)$, we write $X=\Theta(Y)$. We also write $X=o(Y)$ if $|X|\leq c(n)|Y|$ for some function $c(n)\to 0$ as $n\to\infty$. Again, the function $c(\cdot)$ may depend on fixed quantities. We denote by $\|x\|_{{\mathbb{R}}/{\mathbb{Z}}}$ the distance from a real number $x$ to the nearest integer. For $x\in{\mathbb{R}}$, we set $e(x):=e^{2\pi ix}$. For a positive integer $n$, we write $n]=\{1,2,\ldots,n\}$. For a vector ${\bm{v}}\in{\mathbb{R}}^{n}$ and a permutation $\pi$ of $[n]$, we write $\pi({\bm{v}})$ for the vector $(v_{\pi(1)},\dots,v_{\pi(n)})$.
@@
Let $W$ denotes the multi-set $\{w_{1},\dots,w_{n}\}$, and define
@@
Assume that $\rho(W)\geq n^{-C}$ for some constant $C>0$. Then at least $2^{n}n^{-C}$ of the $2^{n}$ subset sums of $W$ coincide, suggesting that $W$ must possess substantial additive structure. Tao and the last author formalized this intuition quantitatively via the notion of generalized arithmetic progressions (GAPs).
@@
[42, Theorem 2.1] Let $C>0$ and ${\varepsilon}\in(0,1)$ be constants. Suppose $W=\{w_{1},\dots,w_{n}\}$ is a multiset of real numbers such that
@@
There are many extensions to other settings, such as the extension from linear forms $S$ of the $\xi_{i}$ to quadratic and higher-order multilinear forms by Costello and Nguyen [10, 39], and Meka et al. [38], as well as extensions to non-abelian groups by Tiep–Vu [47] and Juskevicius–Semetulskis [29] (see also [32, 41, 58]). Other notable contributions include the works of Tao–Vu [62], Rudelson–Vershynin [51], and, more recently, Fox et al. [16] and Kwan et al. [33, 34]. A detailed discussion of these papers is beyond the scope of this work.
The continuous case proceeds in parallel with the discrete case, and we refer the reader to the survey [43] for further discussion. In particular, the continuous analog of Theorem 1.6 asserts that the elements of $W$ lie close to a small GAP.
1.2. New setting: the space of random permutations (discrete setting)
@@
1.2.1. Recent forward theorems
In this subsection, we survey recent results on bounding the probability that $S_{\pi}=x$, for any fixed value $x$. In [54], Söze considered the special case ${\bm{v}}=(1,2,\dots,n)$ and proved the following result.
@@ 1.2.1. Recent forward theorems
$\sup_{x}{\mathbb{P}}(S_{\pi}=x)=O\Big(\frac{1}{n}\Big).$
@@
$\sup_{x}{\mathbb{P}}(S_{\pi}=x)=O\Big(\frac{1}{n}\Big).$
Both Theorems 1.7 and 1.8 are special cases of continuous theorems that we will discuss later. Theorem 1.8 may be viewed as a permutation analog of the Littlewood–Offord–Erdős result from the product-space setting. Note that there is a $\sqrt{n}$ improvement from the bound $O(n^{-1/2})$ to $O(n^{-1})$.
@@
The proof of Theorem 1.9 is quite involved and relies on tools from algebraic combinatorics. In spirit, it is similar to Stanley’s proof [55] of an optimal variant of Theorem 1.2. Roughly speaking, if $\bm{v}$} has distinct coordinates and $S_{\pi}=x$, then the set of such permutations $\pi$ forms an antichain (in a suitable weakening of the Bruhat order on $S_{n}$). This antichain can be identified with a disjoint union of Bruhat orders on $S_{n}/S_{\alpha}$ for certain parabolic subgroups $S_{\alpha}$. These posets have the Sperner property, which implies that their size is bounded by the largest rank of $S_{n}/S_{\alpha}$. These ranks correspond to coefficients of the $q$-multinomial coefficient $\binom{n}{\alpha}_{q}$, and Theorem 1.9 follows from appropriate bounds on those coefficients.
While Theorem 1.9 is elegant—being optimal and valid for all nn—the above methods do not appear to extend easily to more general settings where additional structure is imposed on $\bm{v}$} and $\bm{w}$}. As observed by Alon et al. [2], if both the $v_{i}$ and the $w_{i}$ are distinct, then ${\mathbb{P}}(S_{\pi}=x)$ must lie between $n^{-3/2}$ and $n^{-5/2}$. A sharp bound in this regime would serve as a permutation-space analog of the Erdős–Moser result in the product-space setting. However, the techniques of [2] and [46] do not resolve this problem. We will address it in Section 3.
2. New results in the space of random permutations
We start with highlights of our new results in this paper:
  • In the discrete setting, we prove an analog of Theorem 1.6, namely Theorem 2.1, in the space of random permutations. In fact, we can prove a more general inverse result concerning two-dimensional arrays (matrices), Theorem 2.2. The permutation setting is the special case when the array (matrix) has rank 1.
  • In the continuous setting, we deal with the small-ball probability ${\mathbb{P}}(|S_{\pi}-L|\leq\delta)$. A remarkable feature here (compared to discrete case) is the dependence of the bound on the parameter $L$. Our new results Theorems 2.6, 2.8, and 2.9 establish sub-gaussian behavior with respect to $L$. This is optimal and answers a question posed by Söze [54]. Another key result in this direction, Theorem 2.10, sheds light on the behavior of joint distribution, showing that under certain conditions, two small ball events behave almost independently.
  • As direct applications of Theorem 2.1, we obtain several forward results, such as Theorems 3.1 , 3.2 and $3.$ In particular, we succesfully address the range $n^{-3/2}$ and $n^{-5/2}$ discussed earlier, answering a question of Alon et al. [2].
  • Anti-concentration results in the product setting have a large number of applications in different fields; see [43] for a survey. We expect that our new results in the random permutation setting will have a similar impact. We present two illustrations concerning random polynomials and random matrices.
  • We consider random polynomials of the form $P_{\pi}(x)=\sum_{i=1}^{n}w_{\pi(i)}x^{i-1}$, where ${\bm{w}}=(w_{1},\dots,w_{n})$ is a deterministic real vector and $\pi$ is a random permutation. A well known theorem of Söze [54] shows, under a mild assumption on $\bm{w}$}, that the number of real roots of $P_{\pi}$ is of order $O(\log n)$. We strengthen this result by showing that the number of extremal points (real roots of any fixed order derivative) of $P_{\pi}(x)$ is also of order $O(\log n)$; see Theorem 3.7.
  • Regarding random matrices, we consider the matrix $Q_{n\times n}$ whose rows are $\pi_{1}({\bm{w}}),\dots,\pi_{n}({\bm{w}})$, where $\bm{w}$} is a deterministic vector and $\pi_{1},\dots,\pi_{n}$ are independent random permutations. We show, under a mild assumption on $\bm{w}$}, that $Q_{n\times n}$ is non-singular with high probability; see Theorem 3.9. This theorem can be viewed as the permutation analog of the classical matrix singularity problem in product spaces.
2.1. A new inverse theorem
One of the main results of our paper is the following analog of Theorem 1.6, providing a characterization of those vectors $\bm{w}$} and $\bm{v}$} for which $S_{\pi}$ admits polynomially large concentration.
@@ 2.1. A new inverse theorem
Then for any $n^{{\varepsilon}}\leq n^{\prime}\leq n$ there exists a proper symmetric GAP of rank $r=O_{C,{\varepsilon}}(1)$ and size $O_{{\varepsilon}}(\rho^{-1}(n^{\prime})^{-r/2})$ that contains $(w_{i}-w_{j})(v_{k}-v_{l})$ for at least $(1-n^{\prime}/n)n^{4}$ quadruples $(i,j,k,l)$. In particular, there exists a proper symmetric GAP of rank $r=O_{C,{\varepsilon}}(1)$ and size $O_{{\varepsilon}}(\rho^{-1}n^{-r/2})$ that contains $(w_{i}-w_{j})(v_{k}-v_{l})$ for at least $(1-{\varepsilon})n^{4}$ quadruples.
The bound is sharp; we refer the reader to Lemma 5.1 for a proof. In Section 3, we will use this theorem to derive and refine various forward results. In particular, we will resolve the question posed by Alon et al. discussed in the previous subsection.
2.2. A more general theorem on two dimensional arrays
@@
Then for any $n^{{\varepsilon}}\leq n^{\prime}\leq n$ there exists a proper symmetric GAP of rank $r=O_{C,{\varepsilon}}(1)$ and size $O_{{\varepsilon}}(\rho^{-1}(n^{\prime})^{-r/2})$ that contains $a_{ik}-a_{jk}-a_{il}+a_{jl}$ for at least $(1-n^{\prime}/n)n^{4}$ quadruples $(i,j,k,l)$.
Theorem 2.1 corresponds to the special case $a_{ij}=w_{i}v_{j}$.
2.3. New bounds on the small ball probability: The subgaussian behavior
We assume that $\|{\bm{v}}\|_{\infty}=O(1)$. We inverstigate the small-ball probability ${\mathbb{P}}(|S_{\pi}-L|\leq\delta)$. It is clear that this continuous problem contains the discrete problem considered earlier as a special case with $\delta=0$.
In the product-space setting, this problem has been studied extensively, with early foundational results due to Rogozin [50], Kolmogorov [31], and Halász [18], dated back to the 1960s and 1970s. There have also been many recent developments, especially in the inverse direction; see, for instance, [42, 51, 63] and the survey [43]. For $S_{\pi}$, CLT results and Berry–Esseen–type estimates have been investigated beginning with the classical works of Wald–Wolfowitz [45, 67] and Hoeffding [19] in the early 1950s. More recent approaches based on the Lindeberg exchange method or Stein’s method can be found in [1, 6, 7, 8, 48]. We refer the reader to [9] and the references therein for further contributions concerning this important statistic. The best anti-concentration bound obtainable from these classical approaches is of order $O(n^{-1/2})$ (independent of the length of the interval), arising from the rate of convergence in the CLT.
A key feature of this part is the dependence of the bounds on the parameter $L$. We show that these bounds exhibit sub-gaussian decay in $L$. This behavior is consistent with the central limit theorem and answers a question raised by Söze in [54].
In what follows, we restrict our attention to one of the most natural choices for $\bm{v}$}, namely sequences arising from polynomials. We use this setting in one of our main applications. Our method applies in a more general setting, but the treatment is more technical and will appear elsewhere.
@@ 2.3.1. A starting point: the linear case $v_{i}=i/n$
Our starting point is the following result of Söze [54, Lemma 4], in which he studied $S_{\pi}$ from (2) for the special choice of $\bm{v}$} given by $v_{i}=i/n$.
@@
An interesting feature of this theorem is the appearance of the parameter $L$ in the bound, which implies that $S_{\pi}$ exhibits exponential decay. Söze [54] conjectured a sharper estimate of order $O(n^{-1}e^{-\Theta(L^{2})})$, corresponding to sub-gaussian decay. Notice that once $L$ appears on the right-hand side, the magnitude of the $v_{i}$ becomes relevant.
@@
Other natural questions include weakening the restriction ${\bm{v}}=(1,\dots,n)/n$ and treating scales smaller than $1/n$. We achieve these goals under a modest assumption on $\bm{w}$}. In what follows, we impose the following assumption.
Condition 2.5 (Non-degeneracy).
@@
Roughly speaking, this condition corresponds to $\bm{w}$} being a unit vector orthogonal to $\mathbf{1}$ and satisfying $\|{\bm{w}}\|_{\infty}=o(1/\sqrt{\log n})$.
@@
Let $\delta>0$ be given. Suppose that the sequence $(w_{1},\ldots,w_{n})$ satisfies Condition 2.5 for some sufficiently large constant $A>0$. Let $I\subset[n]$ be any subset with $|I|\geq\delta n$, and consider a sequence $(v_{1},\ldots,v_{n})$ that is partially specified by
@@
${\mathbb{P}}\Bigl(\bigl|S_{\pi}-L\bigr|\leq\frac{1}{n}\Bigr)=O_{A}\Big(\frac{1}{n}\Big).$ (5)
@@
${\mathbb{P}}\Bigl(\bigl|S_{\pi}-L\bigr|\leq\frac{1}{n}\Bigr)=O\Big(\frac{1}{n}e^{-\Theta(L^{2})}\Big).$ (6)
@@
The answer to Problem 2.4 is affirmative for any vector $\bm{w}$} satisfying Condition 2.5.
@@
Let $0<\varepsilon<1/2$ and $\delta>0$ be given. Suppose that the sequence $(w_{1},\ldots,w_{n})$ satisfies Condition 2.5 for some sufficiently large constant $A>0$. Furthermore, assume that no interval of length $\varepsilon/\sqrt{n}$ contains more than $(1-\varepsilon)n$ of the values $w_{i}$. Let $I\subset[n]$ be any subset with $|I|\geq\delta n$, and consider a sequence $(v_{1},\ldots,v_{n})$ that is partially specified by
@@
${\mathbb{P}}\Bigl(\bigl|S_{\pi}-L\bigr|\leq\frac{1}{n^{3/2}}\Bigr)=O_{A}\Big(\frac{1}{n^{3/2}}\Big).$ (7)
@@
${\mathbb{P}}\Bigl(\bigl|S_{\pi}-L\bigr|\leq\frac{1}{n^{3/2-{\varepsilon}}}\Bigr)=O\Big(\frac{1}{n^{3/2-{\varepsilon}}}e^{-\Theta(L^{2})}\Big).$ (8)
@@
This result may be viewed as a continuous analog of Theorem 3.2. We emphasize that the condition excluding intervals of length $1/\sqrt{n}$ that contain most of the $w_{i}$ is essential for obtaining small-ball estimates at the $n^{-3/2}$ scale. In the $L$-dependent bound (8), we assume the slightly larger radius $n^{-3/2+\varepsilon}$ purely for technical convenience; see our treatment of the “very large |t||t|” regime in the proof of (8) in Section 6.
@@
Finally, we remark that Theorem 2.6 may still hold without Condition 2.5. However, for Theorems 2.8 and 6.4, additional assumptions (of the type imposed above) are necessary, since these results fail dramatically when only a few of the $w_{i}$ are nonzero.